one solution is 20% acid and another is 70% acid. How many liters of each solution should you use to make 4 liters of a 55% solution
Answer by mananth(16949)

percent ----------------quantity
Acid type I 20 ----------------x liters
Acid type II 70 ----------------4-x liters
Mixture 55.00% ---------------- 4

20x+70(4-x)=55*4

20x+280-70x=220
20x-70x=220-280
-50x=-60
/-50
x=1.2 liters 20.00% Acid type I
2.8 liters 70.00% Acid type II